OpenStudy (anonymous):

Quadratic Formula..please help. 39=12x-3x^2

OpenStudy (whpalmer4):

Subtract 39 from both sides to arrange in form \[0=ax^2+bx+c\] Then solve with quadratic formula\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

if it was me, i would rewrite \(39=12x-3x^2\) as \[3x^2-12x+39=0\] first then divide all by 3 and get \[x^2-4x+13=0\] and then use teh quadratic formula the zeros, however, will be complex

OpenStudy (whpalmer4):

Yes, factoring out the 3 will make the numbers a bit more arithmetic-friendly. Just so there's no room for confusion, factoring out the 3 doesn't change the answer, and you will get complex numbers either way.